OpenStudy (anonymous):

Can someone help me?? Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

-3|x + 7| = -12 |x+7| = 4 (x+7) = -4 or 4 x = 4-7 or -4-7

OpenStudy (gorv):

|x+7|=4 x+7=-4 or +4 x=-4-7 or +4-7 x=-11 or -3

OpenStudy (gorv):

| | mode symbol use to make negative value positive